OpenStudy (anonymous):

The distance between the two parallel lines L1: −7 x − 10 y=−4, L2: −7 x − 10 y=3 is ???

OpenStudy (anonymous):

its not 7 i've tried that

OpenStudy (anonymous):

its not 7/10 or .7 either

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i got to 7 myself and tried it on the answers it said its wrong

OpenStudy (anonymous):

Its not, graph it!

OpenStudy (anonymous):

And then measure it yourself :-D.

OpenStudy (anonymous):

its not 1 either

OpenStudy (anonymous):

What do the answers say?

OpenStudy (anonymous):

when i punch it in says its incorrect

OpenStudy (anonymous):

Sooory

OpenStudy (anonymous):

its 0.7!

OpenStudy (anonymous):

Because \[ -7x-10y=c \] gives you \[y=-7/10 x - c/10 \] Plug in -4 and 3 and you get 4/10 as opposed to -3/10.

OpenStudy (anonymous):

And 4/10 - (-3/10) = 7/10 = 0.7

myininaya (myininaya):

the answer is 7/10 since we have -7x-10y=-4=> y=-7x/10+4/10 -7x-10y=3=> y=-7x/10-3/10 so they have a distance of 4/10-(-3/10)=7/10 from each other

OpenStudy (anonymous):

it says its incorrect :"(

OpenStudy (anonymous):

But it is not incorrect. That is the right answer, we just proved it.

OpenStudy (anonymous):

Your answer sheet is flawed.

myininaya (myininaya):

did you write down the right equations

OpenStudy (anonymous):

yeah copy paste the same exact thing

myininaya (myininaya):

−7 x − 10 y=−4 −7 x − 10 y=3